1.         How many milligrams would a patient receive if the order is for 15 ml and the         medication strength is 300 mg/tsp?

            Give _____________________

2.         The order: Lanoxin 0.125 mg po every day

            Label:        Lanoxin 0.25 mg/tablet

            Give _____________________

3.         Order: Codeine elixir gr i po every 4 hours as needed for pain (gr = grain)

            Label:   Codeine elixir gr ½ per 5 ml

            Give _____________________

4.         Order: Ancef 40 mg/kg/day in divided doses every 8 hours for infection

            Supply: Ancef 1 gram (1 G)/ 5 mL              Child’s weight: 30 lb.

            How many mL would be given to provide the ordered dose per administration?     

           a.          0.5 mL

           b.          0.7 mL

           c.          0.9 mL

           d.          1.4 mL

5.         An order reads: give Tylenol grains 10 if oral temperature is above 101°F or above.

            The patient’s temperature is 39.2°C.

            What is the patient’s temp in Fahrenheit (F)?   _________________

            Will you give the medication?     YES        NO      (circle one)

           

            Tylenol is available in tablets 1 tab = 325 mg

            If needed, the patient will receive ___________________ tab(s)

6.         Order: Kantrex 15 mg/kg/day IM in four divided doses to an adult weighing 157 lb.

            Label: Kantrex 1gram/3ml

            Give _______________________

7.         Order: 800 ml of D5 L/R in 4 hours

            Drop factor: 15 gtts/ml

            gtts/min________________________

           

            ml/hr___________________________

8.         Order: 50 ml of fluid to infuse in 30 minutes

            Drop factor: 60 gtts/ml

            gtts/min________________________

           

            ml/hr___________________________

9.         Order: Phenobarbitol Sodium 100 mg IM.

            Label: Phenobarbitol Sodium 125 mg / 2 ml.

            Give ______________ml.

10.       Order: Keflin 500 mg IM.

            Label: Keflin 1 g

            Reconstitution: add 4 ml of sterile diluent to yield 0.5 g / 2.2 ml.                    

           

            Give ______________ml.

11.       Compute the total eight hour intake in mL:      

            po: 500 ml water, 2 cups soup, 8 oz juice

            IV: 100 ml per hour

            I = __________________________

12.       Order: Give Tylenol gr XV for temp > 101 degrees F.

            Temp is 101.8 degrees F.    

            Medication on hand is Tylenol tabs: 500 mg = 1 tab.

            (gr XV = 1 gram; gr i = 60 mg)

           How many tabs do we give? ___________________________tab(s)
Answer #1
  1. Order = 15 ml

Available medication strength=300 mg/tsp

                                        = 300mg /5ml (as 1tsp= 5ml)

                                          =60mg /ml

For 15 ml give= 15/60= 0.25 ml

  1. Ordered Lanoxin= 0.125 mg po every day, po means orally

Labelled Lanoxin= 0.25 mg/tablet

              Required tablet= 0.125/0.250= 0.5 tablet means half tablet should be given

3. Ordered codeine elixir gr i po every 4 hours as needed for pain (gr = grain)

           

      Labelled codeine elixir= gr ½ per 5 ml.

                                      1 gr in 10 ml

Hence 10 ml should be givesn every 4 hours for pain

  1. Ordered Ancef = 40 mg/kg/day in divided doses every 8 hours for infection

    

Child’s weight: 30 lb= 30/2.2=13.636 kg

As per weight required dose= 13.636 X 40=545.45 mg

How many mL would be given to provide the ordered dose per administration?

Supplied Ancef =1 gram (1 G)/ 5 mL    

                           = 1000 mg/5 ml

                           = 200mg/ml

For 545.45 mg required ml= 545.45/200== 2.727 ml   per day

Per dose= 2.7/3= 0.9 ml (as every 8 hours means in 24 hours 3 times)

       Hence option C is correct.

5.Order is to give Tylenol grains 10 if oral temperature is above 101°F or above.

The patient’s temperature is 39.2°C.

The patient’s temperature in Fahrenheit (F)? = 102.56

As the temperature is above 101°F hence       YES     

   Tylenol is available in tablets 1 tab = 325 mg  

grain 10= 650 mg

Hence required tablet= 650/325= 0.5 or half tablet should be given.

6.Ordered Kantrex= 15 mg/kg/day IM in four divided doses

Patient’s weight= 157 lb.= 157/2.2= 71.36kg

As per weight dose= 15 X 71.36= 1070.45 mg

Labelled Kantrex =1gram/3ml

                                 = 1000 mg/3ml=333.33 mg/ml

For 1070.45 mg required ml= 1070.45/333.33 ml

                                             = 3.2 ml per day in four divided doses

Per dose= 3.2/4= 0.8 ml

                           

7. Order= 800 ml of D5 L/R in 4 hours

            Drop factor: 15 gtts/ml

          

gtts/min= (800ml/240) x 15= 50 gtts/min

           

            ml/hr= 800/4

                      = 200 ml /hr

8. Order: 50 ml of fluid to infuse in 30 minutes

            Drop factor: 60 gtts/ml

            gtts/min= (50/30) X 60= 100

                         = 100gtts/min

            ml/hr= 50/(30/60)

                   = (50 X60)/30

                        = 100ml/hr

9. Ordered Phenobarbitol Sodium =100 mg IM.

            Labelled Phenobarbitol Sodium= 125 mg / 2 ml.= 62.5 mg/ml

For 100 mg required ml= 100/62.5

                                      = 1.6 ml

10. Ordered Keflin =500 mg IM.

            Label: Keflin 1 g

            Reconstitution: add 4 ml of sterile diluent to yield 0.5 g / 2.2 ml.  

means 500 mg in 2.2 ml

Hence for 500 mg required ml = 2.2 ml